# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v1

@@ 3 Recovery by maximizing the size of intersection graph
Let $\mathcal{T}_{k}\subseteq\mathcal{S}_{n,m}$ denote the set of injections $\pi$ such that $d(\pi,\pi^{*})=k$. For $\pi\in\mathcal{T}_{k}$, by definition, there exists a set of correctly matched vertices (the self-loops in the correlated functional digraph of $\pi$ over the vertices), denoted by $F_{\pi}\triangleq\{v\in S^{*}\cap S:\pi^{*}(v)=\pi(v)\}$ of cardinality $|F_{\pi}|=m-k$. The induced subgraphs of ${\mathcal{H}}_{\pi}$ and ${\mathcal{H}}_{\pi^{*}}$ over $F_{\pi}$ are identical. Therefore,
$e({\mathcal{H}}_{\pi})\geq e({\mathcal{H}}_{\pi^{*}})\iff e({\mathcal{H}}_{\pi})-e({\mathcal{H}}_{\pi}[F_{\pi}])\geq e({\mathcal{H}}_{\pi^{*}})-e({\mathcal{H}}_{\pi^{*}}[F_{\pi}]).$
@@
It should be noted that correlated random variables are contained within the two sides of the inequality. Nevertheless, for any threshold $\tau_{k}$, either $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}$ or $\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}$ holds. Therefore, we have the following upper bound:
$\{d(\hat{\pi},\pi^{*})=k\}\subseteq\bigcup_{\pi\in{\mathcal{T}}_{k}}\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\}\cup\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\}.$
@@
The first event is indicative of a weak signal, while the latter implies the presence of strong noise. The crucial result to establish is that, for a suitable threshold $\tau_{k}$, both bad events will occur with a low probability. Here we may pick $\tau_{k}$ a function of all other parameters $m,k,p,\rho$. For brevity we also write $\tau_{k}=\tau(m,k,p,\rho)$.
@@
For a fixed $\pi\in\mathcal{T}_{k}$, the random variable $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])$ counts the total number of edges among $N_{k}\triangleq\binom{m}{2}-\binom{m-k}{2}=mk(1-\frac{k+1}{2m})$ pairs of vertices, where each edge presents independently with probability $p_{11}$. Furthermore, $F_{\pi}$ is a subset of $S^{*}$ of cardinality $m-k$. While the size of ${\mathcal{T}}_{k}$ could be large, the total number of possible $F_{\pi}$ is at most $\binom{m}{m-k}=\binom{m}{k}$. Therefore,
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\mathbb{P}\left[\bigcup_{\begin{subarray}{c}F\subseteq S^{*}\\ |F|=m-k\end{subarray}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F])<\tau_{k}\right\}\right]$
$\leq\binom{m}{k}\mathbb{P}\left[\mathrm{Bin}(N_{k},p_{11})<\tau_{k}\right].$ (9)
For $\tau_{k}<N_{k}p_{11}$, the tail of binomial distributions follows from the standard Chernoff bound.
@@
The analyses for the noise part is more involved due to the mismatch between $\pi$ and the underlying $\pi^{*}$. Let $S_{\pi}$ denote the domain of $\pi$, and ${\mathcal{E}}_{\pi}\triangleq\binom{S_{\pi}}{2}-\binom{F_{\pi}}{2}$. Then the total number of edges $\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])$ can be equivalently represented as $|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|$, and the cumulant generating function has been upper bounded in Lemma 1 thanks to the decomposition based on the correlated functional digraph. Thus, the error probability can be obtained via the Chernoff bound by optimizing over $t>0$ in (8).
To this end, we need to upper bound the number of self-loops in (8). For a self-loop over an edge $e=uv$, we have $\pi(uv)=\pi^{*}(uv)$. Note that ${\mathcal{E}}_{\pi}$ excludes the edges in the induced subgraph over $F_{\pi}$. It necessarily holds that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$, which contributes two mismatched vertices in the reconstruction of the underlying mapping. Since the total number of mismatched vertices for $\pi\in{\mathcal{T}}_{k}$ equals to $k$, the number of self-loops is at most $\frac{k}{2}$. Consequently, applying (8) with the formula of lower-order cumulants (6) and (7) yields the following lemma, whose proof is deferred to Section A.2.
@@
If $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$, then
$\mathbb{P}\left[|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$
In view of Lemma 2, for $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$, we can apply the union bound for the probability of the bad event due to noise. It remains to upper bound the cardinality of $\mathcal{T}_{k}$. We first choose $m-k$ elements from the domain of $\pi^{*}$ and map them to the same value as $\pi^{*}$. Then, the remaining domain and range of size $k$ and the mapping are selected arbitrarily. Then we obtain
$|\mathcal{T}_{k}|\leq\tbinom{m}{m-k}\tbinom{n-m+k}{k}^{2}k!\leq\frac{m^{k}n^{2k}}{k!^{2}},$
where the last step applies the upper bound $\binom{n}{k}\leq\frac{n^{k}}{k!}$. Since $e^{\frac{k\gamma}{4(2+\gamma)}}\leq e^{k/4}$, we have
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$ $=\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}\right\}\right]$
$\leq|{\mathcal{T}}_{k}|e^{\frac{k\gamma}{4(2+\gamma)}}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$\leq n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right).$ (10)
The following propositions provide sufficient conditions on $m$ for partial and exact recoveries.
Proposition 1 (Upper bound for partial recovery).
@@
For any $\delta\in(0,1)$, there exists a constant $c_{1}(\delta)>0$ such that, when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\leq(\log n)^{-1+o(1)}.$
Proposition 2 (Upper bound for exact recovery).
@@
There exists a universal constant $C>0$ such that, when $m\geq C\left(\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\vee\frac{\log n}{p^{2}\phi(\gamma)}\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{\exp(-\log m)}{1-\exp(-\log m)}+\frac{\exp(-\log n)}{1-\exp(-\log n)}.$
By Propositions 1 and 2, we prove the possibility results in Theorems 1 and 2.
3.1 Proof of Proposition 1
Let $\tau_{k}=N_{k}p_{11}(1-\eta)$ with
$\eta=\sqrt{\frac{8h(\frac{k}{m})}{kp_{11}}}\cdot\mathbbm{1}_{k\leq m-1}+\sqrt{\frac{\log n}{kmp_{11}}}\cdot\mathbbm{1}_{k=m},$
where $h(x)\triangleq-x\log x-(1-x)\log(1-x)$ is the binary entropy function. Since $h(x)/x$ decreases in $(0,1)$, $h(x)/x\geq h(1-\delta)/(1-\delta)$ for $11-\delta\leq x<$ 1. By Lemma 17.5.1 in [CT06], we have $\tbinom{m}{k}\leq\exp\left[mh(k/m)\right]$ for any $k\leq m-1$. By (9) and the Chernoff bound (21), when $k\leq m-1$, we have

### v2

@@ 3 Recovery by maximizing total similarity score
For any $\pi\in\mathcal{T}_{k}$, by definition, there exists a set of correctly matched vertices of cardinality $m-k$—corresponding to the self-loops in the correlated functional digraph of $\pi$ over the vertices—denoted by $F_{\pi}\triangleq\{v\in S^{*}\cap S:\pi^{*}(v)=\pi(v)\}$ with $|F_{\pi}|=m-k$. The induced subgraphs of ${\mathcal{H}}_{\pi}^{f}$ and ${\mathcal{H}}_{\pi^{*}}^{f}$ over $F_{\pi}$ are identical. Consequently,
$\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\iff\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}]).$
@@
It should be noted that correlated random variables are present on both sides of the inequality. Nevertheless, for any threshold $\tau_{k}$, either $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}$ or $\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}$ must hold. This leads to the following upper bound:
$\left\{\max_{\pi\in\mathcal{T}_{k}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\right\}\subseteq\bigcup_{\pi\in{\mathcal{T}}_{k}}\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\}\cup\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\}.$ (10)
@@
The first event indicates the presence of a weak signal, while the second reflects the impact of strong noise. The key result to establish is that, for a suitable $\tau_{k}$, the probabilities of these bad events are small. Here, $\tau_{k}$ can be selected as a function of $m,k,p,\rho$, and $f$. For brevity, we write $\tau_{k}=\tau(m,k,p,\rho,f)$.
@@
For a fixed $\pi\in\mathcal{T}_{k}$, the random variable $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])$ represents the total weight across $N_{k}\triangleq\binom{m}{2}-\binom{m-k}{2}=mk(1-\frac{k+1}{2m})$ pairs of vertices. Additionally, $F_{\pi}$ is a subset of $S^{*}$ of cardinality $m-k$, and the total number of possible configurations for $F_{\pi}$ is at most $\binom{m}{m-k}=\binom{m}{k}$. Therefore,
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\mathbb{P}\left[\bigcup_{\begin{subarray}{c}F\subseteq S^{*}\\ |F|=m-k\end{subarray}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F])<\tau_{k}\right\}\right]$
$\leq\binom{m}{k}\mathbb{P}\left[\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])<\tau_{k}\right].$ (11)
@@
The analysis of the noise component is more challenging due to the mismatch between $\pi$ and the underlying $\pi^{*}$. Let $S_{\pi}$ denote the domain of $\pi$, and ${\mathcal{E}}_{\pi}\triangleq\binom{S_{\pi}}{2}-\binom{F_{\pi}}{2}$ with $|{\mathcal{E}}_{\pi}|=N_{k}$. The total weight $\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])$ can be equivalently expressed as $\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})$. The cumulant generating function for this quantity is upper bounded in Lemma 1, utilizing the decomposition provided by the correlated functional digraph. As a result, the error probability can be evaluated using the Chernoff bound by optimizing over $t>0$ in (8).
To this end, we need to upper bound the number of self-loops in (8). For a self-loop over an edge $e=uv$, we have $\pi(uv)=\pi^{*}(uv)$. Note that ${\mathcal{E}}_{\pi}$ excludes the edges in the induced subgraph over $F_{\pi}$. Therefore, it necessarily holds that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$, which contribute two mismatched vertices in the reconstruction of the underlying mapping. Since the total number of mismatched vertices for $\pi\in{\mathcal{T}}_{k}$ equals $k$, the number of self-loops is at most $\frac{k}{2}$. Consequently, applying (8) with the formula for lower-order cumulants provides an upper bound for $\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]$.
It remains to upper bound the total number of $\pi$ in $\mathcal{T}_{k}$. To do so, we first choose $m-k$ elements from the domain of $\pi^{*}$ and map them to the same value as $\pi^{*}$. Then, the remaining domain and range, each of size $k$, are matched arbitrarily. This yields:
$|\mathcal{T}_{k}|\leq\binom{m}{m-k}\binom{n-m+k}{k}^{2}k!\overset{\mathrm{(a)}}{\leq}\frac{m^{k}n^{2k}}{k!^{2}}\overset{\mathrm{(b)}}{\leq}\frac{n^{3k}}{k!^{2}},$
where $\mathrm{(a)}$ uses the bound $\binom{m}{m-k}\leq\frac{m^{k}}{k!}$ and $\binom{n-m+k}{k}\leq\frac{(n-m+k)^{k}}{k!}\leq\frac{n^{k}}{k!}$, and $\mathrm{(b)}$ is because $m\leq n$. Therefore,
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\right\}\right]\leq\frac{n^{3k}}{k!^{2}}\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right].$ (12)
3.1 Erdős-Rényi model
In this subsection, we focus on the Erdős-Rényi model and use $\hat{\pi}$ from (6) with $f(x,y)=xy$ as the estimator. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])\sim\mathrm{Bin}(N_{k},p_{11})$ with $|F|=k$, the error probability follows from the standard Chernoff bound (see, e.g., (52)): for $10<\eta<$ 1,
$\mathbb{P}\left[\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right]\leq\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right),\quad\tau_{k}=N_{k}p_{11}(1-\eta).$ (13)
For the bad event of noise, applying (8) along with the lower-order cumulant (36) and (37) yields the following lemma, whose proof is deferred to Appendix B.2.
@@ 3.1 Erdős-Rényi model
In the Erdős-Rényi model, for $f(x,y)=xy$ and $\pi\in\mathcal{S}_{n,m}$ with $k=d(\pi,\pi^{*})$, if $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$, then
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$ (14)
Since $p_{11}=p^{2}(1+\gamma)$, we need to choose an appropriate $\eta\in(0,1)$ depending on the model parameters, such that $(1+\gamma)(1-\eta)>1$, in order to apply the error probabilities in (13) and (14) to upper bound (11) and (12), respectively. When $m$ exceeds the sample complexities for partial and exact recovery, we explicitly construct $\eta$ to prove the following propositions, which establish the possibility results stated in Theorems 1 and 2. The proofs of Propositions 1 and 2 are deferred to Appendix A.1 and A.2, respectively.
Proposition 1 (Erdős-Rényi model, upper bound for partial recovery).
@@
For any $\delta\in(0,1)$, there exists a constant $c_{1}(\delta)>0$ such that, when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=xy$ satisfies
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\leq\frac{4\log m+2}{m}.$
Proposition 2 (Erdős-Rényi model, upper bound for exact recovery).
@@
There exists a universal constant $C_{1}>0$ such that, when $m\geq C_{1}\left(\frac{\log(1/p^{2}\gamma)}{p^{2}\gamma}\vee\frac{\log n}{p^{2}\phi(\gamma)}\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=xy$ satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{1}{m-1}+\frac{1}{n-1}.$
3.2 Gaussian Wigner model
In this subsection, we focus on the Gaussian Wigner model. To establish upper bounds for exact recovery, we use different estimators depending on the value of $\rho$, where $0<\rho\leq 1-e^{-12}$ corresponds to the weak signal regime, and $11-e^{-12}<\rho<$ 1 corresponds to the strong signal regime.
Weak signal
Under the weak signal regime $0<\rho\leq 1-e^{-12}$, we use $\hat{\pi}$ in (6) with $f(x,y)=xy$ as our estimator. Similar to Erdős-Rényi model, we upper bound the event $\{d(\hat{\pi},\pi^{*})=k\}$ by the bad event of signal and the bad event of noise according to (10). For the bad event of signal, we note that $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])=\sum_{i=1}^{N_{k}}A_{i}B_{i}$, where $(A_{i},B_{i})$ are independent and identically distributed (i.i.d.) pairs of standard normals with correlation coefficient $\rho$. Then, the error probability can be upper bounded by
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\binom{m}{k}\mathbb{P}\left[\sum_{i=1}^{N_{k}}A_{i}B_{i}<\tau_{k}\right],$ (15)
the tail follows from the Hanson-Wright inequality in Lemma 9. For the bad event of noise, applying (8) with the formula of lower-order cumulants (41) and (42) yields the following Lemma, whose proof is deferred to Appendix B.3.
Lemma 3.
In the Gaussian Wigner model, for $f(x,y)=xy$, any threshold $\tau_{k}$, and $\pi\in\mathcal{S}_{n,m}$ with $k=d(\pi,\pi^{*})$, we have
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{6}+\frac{\rho^{2}|\mathcal{E}_{\pi}|}{14}+\frac{\log 5}{8}k\right).$ (16)
The following proposition provides sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $0<\rho\leq 1-e^{-12}$, whose proof is deferred to Appendix A.3.
Proposition 3 (Gaussian Wigner model, upper bound for exact recovery).
When $0<\rho\leq 1-e^{-12}$, there exists a universal constant $C_{3}>0$ such that, when $m\geq\frac{C_{3}\log n}{\rho^{2}}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=xy$ satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{1}{m-1}+\frac{1}{n-1}.$
Strong signal
Under the strong signal regime $11-e^{-12}<\rho<$ 1, we use $\hat{\pi}$ in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ as our estimator. Indeed, the MLE is also of the form (6), but with $f(x,y)=\rho xy-\frac{\rho^{2}}{2}(x^{2}+y^{2})$, which is approximately $-\frac{1}{2}(x-y)^{2}$ when $\rho=1-o(1)$. This choice reflects the phase transition phenomenon in the Gaussian Wigner model when $\rho=1-o(1)$. Specifically, for a correlated edge pair, the expected mean squared difference $\mathbb{E}\left[\left(\beta_{e}(G_{1})-\beta_{\pi^{*}(e)}(G_{2})\right)^{2}\right]=o(1)$, while for an uncorrelated edge pair, it stays bounded away from zero with high probability. Our proof follows a similar structure with the weak signal regime. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])=\sum_{i=1}^{N_{k}}-\frac{1}{2}(A_{i}-B_{i})^{2}$, where $(A_{i},B_{i})$ are i.i.d. pairs of standard normals with correlation coefficient $\rho$. Then, the error probability can be upper bounded by